Circular No. 7394

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 2000au IN MCG +8-15-14 S. Jha, P. Challis, and R. Kirshner, Harvard-Smithsonian Center for

Astrophysics, report that spectra of SN 2000au (cf. IAUC 7389), taken by W. Brown on Mar. 31.25 and Apr. 3.13 UT with the F. L. Whipple Observatory 1.5-m telescope (+ FAST spectrograph), show it to be a type-II supernova after maximum light, with broad, well-developed Balmer lines and Fe II. Superimposed narrow Hα emission yields a recession velocity of 5900 km/s for the host galaxy. The Hα P-Cyg profile yields an unusually high expansion velocity for the supernova of 10 900 km/s.

SUPERNOVA 2000bg IN NGC 6240 Jha et al. also report that spectra of SN 2000bg (cf. IAUC 7392), taken

as above by Brown on Apr. 3.51 and 4.48 UT, show it to be a type-IIn supernova at an early epoch. The spectra exhibit a blue continuum with a broad Hα emission profile (FWHM = 10 000 km/s) as well as narrow (FWHM < 300 km/s) Hα and [N II] emission. The narrow emission lines yield a recession velocity of 7180 km/s for the host galaxy. The spectra also show narrow interstellar Na I D absorption (EW = 0.50 nm), at the host-galaxy redshift, indicating significant dust absorption.

XTE J1118+480 S. Chaty and C. A. Haswell, Open University; G. P. Smith and I.

Smail, Durham University; and R. I. Hynes, Southampton University, report the following near-infrared photometry of the counterpart to the x-ray transient XTE J1118+480 (IAUC 7389), obtained with the 3.8-m U.K. Infrared Telescope (+ UFTI instrument): Apr. 4.2 UT, J = 12.12 ± 0.02, H = 11.75 ± 0.02, K = 11.06 ± 0.02.

J. Wren, Los Alamos National Laboratory (LANL); and T. McKay, University of Michigan (UM), on behalf of the ROTSE collaboration (UM/ LANL/Lawrence Livermore National Laboratory), report the following magnitudes for the optical counterpart of XTE J1118+480, obtained with the ROTSE-I telescope; all observations were performed using an unfiltered CCD, and magnitudes are calibrated to approximately the V -band (with an uncertainty of 0.1 mag): 1999 Dec. 18.315 UT, [15.3; 28.283, [14.6; 2000 Jan. 4.223, 13.33 ± 0.08; 12.194, 13.10 ± 0.02; 31.143, 14.30 ± 0.08; Feb. 3.147, 14.35 ± 0.08; 8.150, 15.10 ± 0.18; 15.312, [15.0; 23.185, [14.8; 26.299, 14.82 ± 0.12; 29.265, 13.79 ± 0.04; Mar. 5.322, 13.33 ± 0.03; 12.325, 13.14 ± 0.02; 28.206, 13.11 ± 0.03.

2000 April 5  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2000 CBAT  Daniel W. E. Green